Grade 11IIT JEE Entrance Exam

If s=2tcube+3tsquare+2t+8 then the time at which acceleration is zero is

Arpana
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
Dear student
 
s = 2 t³ + 3t² + 2t +8
Now for acceleration, we have to find d²s/dt²
Hence
v = ds/dt = 6t² + 6t + 2
a = dv/dt = d²s/dt² = 12 t +6
Now put a = 0
12t +6 = 0
Hence
t = -1/2
Since t is in negative sign
Hence acceleration can never be zero